OpenStudy (anonymous):

can someone help me with this question. what is the solution set of x^2+11x+n is a perfect square trinomial

OpenStudy (rational):

you want to change it to form : \(\large (x+t)^2\)

OpenStudy (rational):

Notice that you can convert below to a perfect square : \(\large x^2 + 2\times \frac{11}{2}\times x + (\frac{11}{2})^2\)

OpenStudy (anonymous):

ok

OpenStudy (rational):

so \(\large n\) must equal \(\large (\frac{11}{2})^2\) for the given expression to be a perfect square

OpenStudy (anonymous):

so its 30.25

OpenStudy (rational):

yes

OpenStudy (anonymous):

thanks